# DIVERSE: Deciphering Internet Views on the U.S. Military Through Video Comment Stance Analysis, A Novel Benchmark Dataset for Stance Classification

### v1

@@ Abstract
Stance detection of social media text is a key component of downstream tasks involving the identification of groups of users with opposing opinions on contested topics such as vaccination and within arguments. In particular, stance provides an indication of an opinion towards an entity. This paper introduces DIVERSE [^1] The dataset is released at https://doi.org/10.5281/zenodo.10493803, a dataset of over 173,000 YouTube video comments annotated for their stance towards videos of the U.S. military. The stance is annotated through a human-guided, machine-assisted labeling methodology that makes use of weak signals of tone within the sentence as supporting indicators, as opposed to using manual annotations by humans. These weak signals consist of the presence of hate speech and sarcasm, the presence of specific keywords, the sentiment of the text, and the stance inference from two Large Language Models. The weak signals are then consolidated using a data programming model before each comment is annotated with a final stance label. On average, the videos have 200 comments each, and the stance of the comments skews slightly towards the “against” characterization for both the U.S. Army and the videos posted on the channel.
Introduction
At the time of the publication of this article, the United States is facing a military recruitment crisis. For the last few years, the U.S. Military, and the U.S. Army in particular, have struggled to meet recruiting targets to maintain their desired force structure (Brown 2023; Roxana Tiron 2022). Because of the salience of this crisis to the United States’ national well-being (i.e., a professional and ready military is an important component of a healthy state), there has been increased focus from various sectors of government, industry, and academia on discovering why this recruiting crisis is happening and what might be done to address its causes. While there is no agreed-upon reason or set of reasons for why the crisis exists, it is clear that social media plays an important role in military recruiting (Asch 2019; Ng and Cruickshank 2023; Calkins and Asch 2022; Myers 2022; Brown 2023). Thus, it is critical to understand how recruitment messaging is received on social media and how it might be shaping the beliefs of social media users toward something like joining the military. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, whether it be for service in the military or for influencing consumer behavior.
@@ Introduction
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social, social media, and marketing studies. As such, several articles have proposed various methods and datasets for the task of stance detection (Ng and Carley 2022; Allaway and McKeown 2023; Du Bois 2007). However, all existing datasets explore stance as expressed on a single social media platform, Twitter, and only have a single stance per comment (Sobhani, Inkpen, and Zhu 2017; Conforti et al. 2020; Mohammad et al. 2016; Hossain et al. 2020; Kochkina, Liakata, and Zubiaga 2018; Villa-Cox et al. 2020). Furthermore, since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), as well as varying definitions of stance (Ng and Carley 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
To address these shortcomings in prior stance datasets and to aid in understanding the critical issues with the U.S. Army’s recruitment effort, we propose a new stance dataset based on YouTube comments and labeled by human-assisted machine labeling. Our contributions are as follows:
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. These differences include using comments from a different social media platform (YouTube rather than Twitter), comments having labeled stances toward more than one target (the only other being Election2016 (Sobhani, Inkpen, and Zhu 2017)), the presence of contentious issues, conspiracy theories, and other emotional content, and having more than three times the number of labeled comments over the next largest benchmark stance dataset (the next largest benchmark is wtwt at ${}32,000$ data points (Conforti et al. 2020)).
  • The use of a human-guided, machine-assisted stance-labeling methodology. We adopt recent advances in the creation of large, high-quality labeled datasets (Ratner et al. 2016, 2019), combined with knowledge about the domain and stance, to produce a methodology that can generate quality stance labels at a much larger scale than human-annotation-only methods. This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs) to determine the final stance label.
@@
This dataset is a collection of YouTube comments posted on the videos uploaded by the U.S. Army. This set of comments is annotated for weak labels representing tone (hate speech, sarcasm, and sentiment) before combining with predetermined keywords and the inference of two LLMs to describe the final stance of the comment towards the U.S. Army.
Related Work
@@
The task of stance detection has led the community to develop several datasets. Ng and Carley (2022) consolidated several stance detection datasets with social media posts across topics like coronavirus misconceptions, political reactions and tweets revolving around crisis events. Generally, these datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al. 2016). There are datasets that use labels of implicit and explicit support and denial as stances to indicate the overtness of stance expression (Qazvinian et al. 2011). Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab 2016), logistic regressions (Augenstein, Vlachos, and Bontcheva 2016), neural network models (Fang et al. 2019; Siddiqua, Chy, and Aono 2019) and deep learning models (Küçük and Can 2020). Recently, Large-Language Models have also been used for identifying stance in sentences through direct inference and through (Zhang et al. 2023a; Gatto, Sharif, and Preum 2023).
@@
To date, there have been several research investigations involving YouTube, but few datasets. For example, there are a number of papers investigating influence campaigns (Marcoux et al. 2021; Hussain et al. 2018), conspiracy theories (Liaw et al. 2023), comment toxicity (Obadimu et al. 2019), and even cross-platform use of YouTube videos (Ginossar et al. 2022). Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al. 2023). Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection or classification, particularly concerning the comments section on YouTube videos. It is important to note, however, that there are a number of commercially available social media suites that analyze comments to social media posts on sites like YouTube for things like sentiment and influencer engagement, but none of these tools provide stance detection or classification (Ng and Carley 2022; Rogan et al. 2022).
The Use of Large Language Models for Stance Detection
@@
Recently, researchers have explored the application of Large Language Models (LLMs) in stance detection. Existing studies have primarily focused on ChatGPT, yielding mixed outcomes. For example, Zhang et al. demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts, outperforming supervised models (Zhang, Ding, and Jing 2022). Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al. 2023; Liyanage, Gokani, and Mago 2023; Zhang et al. 2023b). Conversely, Aiyappa et al. raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations (Aiyappa et al. 2023). Mets et al. examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models (Mets et al. 2023). Despite these findings, the effectiveness of LLMs in stance classification remains promising, if uncertain, especially without fine-tuning on labeled data or prompt engineering.
Data Programming for Dataset creation
@@
Recently, Data Programming has emerged as a promising approach for expeditiously generating large training datasets. This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al. 2016, 2019). Rather than relying solely on manually annotated datasets, Data Programming enables the automatic creation of training sets by combining various weak sources, such as heuristics, distant supervision, and domain expertise. These weaker, noisy labels can then be combined to produce high-quality labels. This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is prohibitively time-consuming or expensive, such as for stance detection. The incorporation of Data Programming into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
Data
In this section, we detail how the dataset was collected, we describe the annotation process, and we comment on the FAIR aspects of the dataset. For this dataset, we labeled each comment for two stances: the stance of the comment toward the video it is commenting on and the stance of the comment toward the U.S. Army. For each stance, we allowed for three possible options: “supports”, “against”, and “neutral”.
Data Collection
@@
This dataset was collected to understand the reception of YouTube users toward the U.S. Army’s marketing material on YouTube. As mentioned in the introduction, social media marketing is a critical aspect of military recruiting, and military recruiting in the United States is undergoing a crisis. Since the U.S. Army maintains an official channel on YouTube, we collected the comments on the videos of that channel to understand how content from the Army is being received and hopefully gain some insight into the causes behind the recruiting crisis. To do this, we first collected comments from all of the videos posted on the U.S. Army’s official YouTube channel (username: @usarmy [^2] https://www.youtube.com/usarmy) using the YouTube Data API [^3] https://developers.google.com/youtube/v3/docs/comments/list. The collection of comments was only from those videos that enabled comments and were accessible via the YouTube API. We collected both comments as well as any replies to comments. Collection of the comments took place from 2 October 2023 to 5 October 2023. In total, we collected 173,013 comments across 858 videos. For each comment, we collected the following information:
@@
Data Annotation Methodology
@@
To annotate the data for stance labels, instead of using human annotators, we followed the Data Programming Paradigm (Ratner et al. 2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension (Liu and Jaidka 2023) and negotiation strategies using the presence of reasoning and friendliness (Jaidka, Ahuja, and Ng 2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a Data Programming model to produce the final labels. Figure 1, depicts the entire data annotation methodology workflow. In the following subsections, we detail key aspects of the workflow, with a special emphasis on the creation of weak labels.
@@
Figure 1: Pipeline of Stance Dataset Creation Methodology
Pre-tuning of Large Language Models
@@
Based on previous research that has shown the possible utility of stance labeling by LLMs, we opted to use LLMs to label comments for their stances. Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used two different LLMs: Flan-UL2 (Tay et al. 2022), which is an encoder-decoder architecture and Mistral-7B (Jiang et al. 2023), which is a decoder-only architecture. Based on past research (Ziems et al. 2023), we used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model). For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al. 2016), Election2016 (Sobhani, Inkpen, and Zhu 2017), srq (Villa-Cox et al. 2020), wtwt (Conforti et al. 2020), phemerumors (Kochkina, Liakata, and Zubiaga 2018), and covid-lies (Hossain et al. 2020). Based on the research in (Ng and Carley 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al. 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
@@
Weak Labels
To label the comments for stance, we first needed to produce noisy, weak labels. To do this we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. Finally, it is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion (Fortuna and Nunes 2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the U.S. Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection (Fortuna and Nunes 2018). For this weak label, we used the dehatebert model from HuggingFace [^4] Hate-speech-CNERG/dehatebert-mono-english (Aluru et al. 2020), a multi-lingual fine-tuned BERT model. This model was used as is without any fine-tuning. For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the ”hate” response as an “against” stance label.
Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner (Chaudhari and Chandankhede 2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector555helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic (Ibrohim, Bosco, and Basile 2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model666lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
Use of Keywords
Stance from Large Language Models
Based on the recent work demonstrating some effectiveness of LLMs in the stance classification task, we also directly labeled the comments for their stances using LLMs. We used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt. The labeling prompt without incorporating the reply structure is:
@@
As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt. These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt. Since we are labeling the comment for its stance toward the U.S. Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
Use of Other Stances

### v2

# DIVERSE: A Dataset of YouTube Video Comment Stances with a Data Programming Model

@@ Abstract
Stance detection of social media text is a key component of many real-world applications like evaluating marketing campaigns, evaluating political policies or candidates, or evaluating information environments. However, creating automatic stance labeling systems requires the manual annotation of stances, which is both tedious and resource-intensive. This paper introduces an stance labeling method that makes use of weak signals of sentence tone, then consolidating these signals with a Data Programmingmodel for the final stance label. In a time of international conflict, understanding the public opinion towards the country’s military is crucial for recruitment. We present DIVERSE, a dataset involve stances towards YouTube videos of the US military [^1] Datasetavailableathttps://doi.org/10.5281/zenodo.10493803. On average, the videos have 200 comments each, and the stances skew slightly towards the “against” characterization for both the US army and the video.
Index Terms:
stance detection, social media, large language models, ensemble learning
I Introduction
Today’s world is rife with armed conflict: the Russia-Ukraine war, the Israel-Hamas conflict, the Yemeni Civil War, Somali Civil War… and so forth. The US is involved in a number of these conflicts, and also has the mammoth task of defending her own. During conflict, it is crucial to know the stance of the people towards its own defending forces, as a gauge of the support the forces have from the public. However, US military is facing a recruitment crisis. For the last few years, the US Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023). Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023). Thus, it is critical to understand how messages from the military is received on social media, for it may be used to shape beliefs of users. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
@@ I Introduction
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social studies. Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007). However, existing datasets explore stance as expressed on a single social media platform, mainly X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020). Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
@@
To address these shortcomings in prior stance dataset, we propose a new stance dataset based on YouTube comments and labeled with machine labeling through the data programming paradigm of weak supervision. This dataset is a collection of YouTube comments posted on videos uploaded by the US Army [^2] Offical US Army YouTube Channel: https://www.youtube.com/@usarmy. The comments are annotated for weak labels representing tone (hate speech, sarcasm, sentiment), inference of two LLMs (stance towards video, stance towards army), then combining these labels to produce the final stance labels of the comment towards the US Army. With this dataset, we develop a automated and systematic approach to labeling stances, aiding in understanding critical issues with the public opinion towards the army and its recruitment efforts.
Our contributions of this paper are as follows:
  • The use of an data-programming based stance-labeling methodology. We combine recent advances for creating large, high-quality labeled datasets Ratner et al. (2016, 2019) with contextual knowledge to generate quality stance labels. This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. While most stance datasets use posts from Twitter, ours provides posts from the YouTube platform instead, contributing to the breadth of datasets. Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al. (2017)), and provides labels to weak signals including: the presence of contentious issues, conspiracy theories, and other emotional content. Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al. (2020) at ${}32,000$ data points vs 173,013 data points).
II Related Work
@@
The task of stance detection has led to the development of several datasets. Ng and Carley (2022) consolidated several stance detection datasets with social media posts across multiple topics. These datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al., 2016). There are datasets that use labels of implicit and explicit support and denial to indicate the overtness of stance expression (Qazvinian et al., 2011). Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab, 2016), logistic regressions (Augenstein et al., 2016), neural network models (Fang et al., 2019; Siddiqua et al., 2019) and deep learning models (Küçük and Can, 2020). Recently, Large-Language Models have also been used for identifying stance in sentences (Zhang et al., 2023a; Gatto et al., 2023).
@@
To date, there have been several research investigations involving YouTube, but few datasets. For example, there are a number of papers investigating influence campaigns (Marcoux et al., 2021; Hussain et al., 2018), conspiracy theories (Liaw et al., 2023), comment toxicity (Obadimu et al., 2019), and even cross-platform use of YouTube videos (Ginossar et al., 2022). Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al., 2023). Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection, particularly concerning the comments section on YouTube videos. It is important to note, however, that there are a number of commercially available social media suites to analyze comments to posts on platforms for indicators like sentiment and influencer engagement, but none of these tools provide stance detection (Ng and Carley, 2022; Rogan et al., 2022).
Large Language Models for Stance Detection
@@
Recently, researchers have explored the application of Large Language Models (LLMs) in stance detection. Existing studies primarily focused on ChatGPT, yielding mixed outcomes. Zhang et al. (2022) demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts. Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al., 2023; Liyanage et al., 2023; Zhang et al., 2023b). Conversely, Aiyappa et al. (2023) raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations. Mets et al. (2023) examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models. Despite these findings, the effectiveness of LLMs in stance classification remains promising, if uncertain, especially without fine-tuning on labeled data or prompt engineering.
Weak Supervision for Dataset creation
@@
Weak Supervision has emerged as a promising approach for expeditiously generating large training datasets. This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al., 2016, 2019). Rather than relying solely on manually annotated datasets, weak supervision enables the automatic creation of training sets by combining various weak sources, derived from heuristics, distant supervision, and domain expertise. These weaker, noisy labels can then be combined to produce high-quality labels. This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is time-consuming or expensive. The incorporation of weak supervision into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
III Dataset Creation
This dataset contains comments of YouTube comments, each labeled for a set of two final stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army. There are three possible options for stances: “supports”, “against”, and “neutral”. This dataset also contains the weak labels used to fold into the final stance through weak supervision.
III-A Data Collection
@@
This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military. We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list. This returned both comments as well as replies to comments. Collection took place from 2 to 5 October 2023, where we retrieved a total of 173,013 comments across 858 videos. For each comment, we collected the following information:
@@
III-B Data Annotation Methodology
@@
Instead of using human annotators for stance labeling, we used weak supervision Ratner et al. (2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension Liu and Jaidka (2023) and negotiation strategies using the presence of reasoning and friendliness Jaidka et al. (2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a weak supervision model to produce the final labels. Figure 1, depicts the entire data annotation methodology workflow. We detail key aspects of the workflow in the following sections.
@@
Figure 1: Pipeline of stance labeling methodology note: change LLM box to 2 boxes not 3
III-C Annotating Weak Labels
To label the comments for stance, we first produced weak labels. To do this we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. It is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
III-C1 Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion Fortuna and Nunes (2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the US Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018). For this weak label, we used the dehatebert model from HuggingFace [^5] Hate-speech-CNERG/dehatebert-mono-english Aluru et al. (2020), a multi-lingual fine-tuned BERT model. This model was used as is without any fine-tuning. For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the “hate” response as an “against” stance label.
III-C2 Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner Chaudhari and Chandankhede (2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
III-C3 Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic Ibrohim et al. (2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model777lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
III-C4 Use of Keywords
III-C5 Stance from Large Language Models
Based on the recent work demonstrating effectiveness of LLMs in the stance classification task (Ng et al., 2024), we also directly labeled stances using LLMs.
Pre-tuning of LLMs
@@ III-C5 Stance from Large Language Models
Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used two different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture and (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture. We used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023). For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020). Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al., 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
@@
Analyzing stances with LLMs
In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt. The labeling prompt without incorporating the reply structure is:
@@
As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt. These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt. Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
III-C6 Use of Analogous Stances